# Monte Carlo Temperature: a robust sampling strategy for LLM’s uncertainty quantification methods

### v1

@@ 1 Introduction
Large Language Models (LLMs) have fundamentally transformed the way we interact with artificial intelligence, revolutionizing various domains, from content creation to complex problem-solving tasks (Bommasani et al., 2021; Wei et al., 2022; Orrù et al., 2023). However, these powerful models can sometimes produce unreliable or incorrect outputs, raising concerns about their deployment in critical applications (Rohrbach et al., 2018; Xiao & Wang, 2021; Bacciu et al., 2024). While significant research efforts have focused on improving LLMs’ accuracy through techniques like Chain-of-Thought prompting (Wei et al., 2022) and Retrieval-Augmented Generation (Lewis et al., 2020), parallel work has emerged on developing uncertainty quantification (UQ) methods to estimate model confidence as an indicator of potential errors (Kadavath et al., 2022; Kuhn et al., 2023; Lin et al., 2024).
@@
Existing UQ methods for LLMs can be used to predict the correctness of a LLM’s output, either under white-blox or black-box assumptions. They fall into two broad categories: single-sample and multi-sample approaches. Single-sample methods analyze a single generation using metrics like perplexity or evaluating model’s weight activations. In contrast, multi-sample methods, which we focus on in this work, rely on querying the model multiple times with the same input and non-zero fixed temperature sampling, to induce and measure diversity in the generations. To assess the effectiveness of UQ methods in distinguishing between correct and incorrect model outputs, they are typically evaluated as a classification procedure using the area under the receiver operator characteristic curve (AUROC) metric (Hanley & McNeil, 1982). However, the impact of selecting a specific fixed temperature parameter is understudied, and our analysis reveals that temperature plays a fundamental role in the effectiveness of different UQ methods across scenarios in which different LLMs are employed to solve different tasks. Figure 1 exemplifies this behavior over four question-answering datsets and three models using the semantic entropy method [^1] Similar plots for other UQ methods can be found in the Appendix A. (Kuhn et al., 2023). The figure highlights three critical observations: (1) for a given model and dataset, performance varies significantly with changes in temperature; (2) no single temperature consistently optimizes performance across datasets for a given model; and (3) no universal temperature yields optimal results across models for a given dataset. For instance, the Falcon-40B model achieves peak performance on the TriviaQA dataset at a temperature of $0.6$, but requires a lower temperature of $0.3$ for the SVAMP dataset. Similarly, within the same TriviaQA dataset, optimal temperature values differ across different models: Falcon-40B performs best at $0.6$, while Falcon-7B-Instruct achieves superior results at $1.0$. This lack of robustness in maintaining consistent performance across different scenarios poses significant challenges for practitioners attempting to implement UQ methods and highlight the need for more robust approaches to temperature selection.
@@ 3.1 Robustness Definition in UQ Methods
Robustness in the context of UQ refers to the stability and generalizability of a UQ method’s performance when applied across different settings. In our use case, robustness captures the range to which a UQ method remains effective in assessing uncertainty under changes in the following dimensions:
@@ 4.3 Entailment and Evaluation Model
To assess the effectiveness of UQ methods, we measured performance using AUROC, PR-AUC, and AURAC metrics (Hanley & McNeil, 1982; Davis & Goadrich, 2006; Farquhar et al., 2024). Confidence intervals at the 95% level were computed for all metrics via bootstrapping to ensure statistical relevance.
@@ 6 Conclusion
In this work, we introduced MCT, a general and robust sampling method for UQ in LLMs. Our approach eliminates the need for expensive HPO of temperature parameters, providing consistent performance across a wide range of models, datasets, and UQ methods. The experimental results demonstrate that MCT achieves statistical parity with oracle-fixed temperatures obtained through computationally intensive optimization. Additionally, it outperforms the Best On Average and Fixed Random Temperature baselines by reducing performance variability and enhancing robustness across diverse configurations.
@@
Despite its positive results, this study has some limitations. While we validated MCT across a diverse set of UQ techniques and LLMs, further exploration is necessary to evaluate its effectiveness on larger-scale models and alternative architectures. Additionally, this work primarily focused on temperature as the inference parameter; future studies should investigate the impact of other sampling techniques and inference configurations, such as top-P and top-k sampling, to extend MCT’s applicability.

### v2

@@ Abstract
Monte Carlo Temperature: a robust sampling strategy for LLM’s uncertainty quantification methods
Nicola Cecerethanks: Equal contribution,  Andrea Bacciu11footnotemark: 1,  Ignacio Fernández Tobías,  Amin Mantrach Amazon nicola.cecere@mail.polimi.it {andbac, tobiasi, mantrach}@amazon.com
@@ 1 Introduction
Large Language Models (LLMs) have fundamentally transformed the way we interact with artificial intelligence, revolutionizing various domains, from content creation to complex problem-solving tasks (Bommasani et al., 2021; Wei et al., 2022; Orrù et al., 2023). However, these powerful models can sometimes produce unreliable or incorrect outputs, raising concerns about their deployment in critical applications (Rohrbach et al., 2018; Xiao and Wang, 2021; Bacciu et al., 2024). While significant research efforts have focused on improving LLMs’ accuracy through techniques like Chain-of-Thought prompting (Wei et al., 2022) and Retrieval-Augmented Generation (Lewis et al., 2020), parallel work has emerged on developing uncertainty quantification (UQ) methods to estimate model confidence as an indicator of potential errors (Kadavath et al., 2022; Kuhn et al., 2023; Lin et al., 2024).
@@
Existing UQ methods for LLMs can be used to predict the correctness of a LLM’s output, either under white-blox or black-box assumptions. They fall into two broad categories: single-sample and multi-sample approaches. Single-sample methods analyze a single generation using metrics like perplexity or evaluating model’s weight activations. In contrast, multi-sample methods, which we focus on in this work, rely on querying the model multiple times with the same input and non-zero fixed temperature sampling, to induce and measure diversity in the generations. To assess the effectiveness of UQ methods in distinguishing between correct and incorrect model outputs, they are typically evaluated as a classification procedure using the area under the receiver operator characteristic curve (AUROC) metric (Hanley and McNeil, 1982). However, the impact of selecting a specific fixed temperature parameter is understudied, and our analysis reveals that temperature plays a fundamental role in the effectiveness of different UQ methods across scenarios in which different LLMs are employed to solve different tasks. Figure 1 exemplifies this behavior over four question-answering datsets and three models using the semantic entropy method [^1] Similar plots for other UQ methods can be found in the Appendix A. (Kuhn et al., 2023). The figure highlights three critical observations: (1) for a given model and dataset, performance varies significantly with changes in temperature; (2) no single temperature consistently optimizes performance across datasets for a given model; and (3) no universal temperature yields optimal results across models for a given dataset. For instance, the Falcon-40B model achieves peak performance on the TriviaQA dataset at a temperature of $0.6$, but requires a lower temperature of $0.3$ for the SVAMP dataset. Similarly, within the same TriviaQA dataset, optimal temperature values differ across different models: Falcon-40B performs best at $0.6$, while Falcon-7B-Instruct achieves superior results at $1.0$. This lack of robustness in maintaining consistent performance across different scenarios poses significant challenges for practitioners attempting to implement UQ methods and highlight the need for more robust approaches to temperature selection.
@@ 3.1 Robustness Definition in UQ Methods
Robustness in the context of UQ refers to the stability and generalization of a UQ method’s performance when applied across different settings. In our use case, robustness captures the range to which a UQ method remains effective in assessing uncertainty under changes in the following dimensions:
@@ 4.1 Configuration for Generating Answers
Once the parameters $\tau_{\text{min}}$, $\tau_{\text{max}}$, and $k$ are defined, applying equation (3) yields the exact interval that we employed for MCT sampling: $\{0.100,0.325,0.550,0.775,1.000\}$.
@@ 4.3 Entailment and Evaluation Model
To assess the effectiveness of UQ methods, we measured performance using AUROC, PR-AUC, and AURAC metrics (Hanley and McNeil, 1982; Davis and Goadrich, 2006; Farquhar et al., 2024). Confidence intervals at the 95% level were computed for all metrics via bootstrapping to ensure statistical relevance.
@@ 5.2 Comparison with the Best On Average and Fixed Random Temperature
Moreover, this advantage translates into strong win-rate performance for MCT. It outperforms the Best Average Fixed Temperature method in 63.24% of cases and achieves an even greater win rate of 72.03% against the Random Baseline, further confirming its robustness.
@@ 6 Conclusion
In this work, we introduced MCT, a general and robust sampling method for UQ in LLMs. Our approach eliminates the need for expensive HPO of temperature parameters, providing consistent performance across a wide range of models, datasets, and UQ methods. The experimental results demonstrate that MCT achieves statistical parity with oracle-fixed temperatures obtained through computationally intensive optimization. Additionally, it outperforms the best average fixed-temperature and random baselines by reducing performance variability and enhancing robustness across diverse configurations.
@@
7 Limitations
While this study demonstrates promising results, several limitations must be acknowledged. Although we validated MCT across a diverse set of UQ techniques and LLMs, further exploration is required to assess its effectiveness on larger-scale models and alternative architectures. Additionally, this work primarily focused on temperature as the inference parameter; future studies should examine the impact of other sampling techniques and inference configurations, such as top-P and top-k sampling, to expand MCT’s applicability.
8 Acknowledge
We want to thank Marcello Federico for his valuable support and feedback on this paper.